OpenStudy (caerus):

Guys can you help me with this..Linear equation of order one, i try and cant do it

OpenStudy (caerus):

\[x(x^2+1)y'+2y= (x^2+1)^3\]

OpenStudy (caerus):

i will post what i done and correct me with my error ^.^

OpenStudy (jiteshmeghwal9):

\[x(x^2+1)y'+2y=(x^2+1)^3\]\[y'+\frac{2}{x(x^2+1)}y=\frac{(x^2+1)^3}{x(x^2+1)}\]

OpenStudy (jiteshmeghwal9):

form of a linear differential equation is \[y'+Py=Q\]\[\text{Integrating factor (I.F.)}=e^{\int{Pdx}}\] final solution of the equation =\[y* \text{I.F.}= \int\limits Q* \text{I.F.} dx +k\]

OpenStudy (caerus):

\[p \int\limits_{}^{}\frac{ 2 }{ x(x^2+1) }dx\]

OpenStudy (jiteshmeghwal9):

In ur question \[Q= \frac{(x^2+1)^3}{x(x^2+1)}\]\[\Large \text{I.F.}=e^{\int {\frac{2}{x(x^2+1)}dx}}\]

OpenStudy (caerus):

how can i integrate that :)

OpenStudy (caerus):

i will distribute the denominator? \[\int\limits_{}^{} \frac{ 2}{ x^3+x }dx\]

OpenStudy (caerus):

\[2((x^-2/-2)+(lnx))\]

OpenStudy (caerus):

\[e^{-x^-2}+x+c\]

OpenStudy (caerus):

i end up \[x e^{-x^-2}\]

OpenStudy (caerus):

\[\int\limits_{}^{}Q(x).I.F\]

OpenStudy (caerus):

\[yxe ^{-x^-2}=\int\limits_{}^{}\frac{ (x^2+1)^2 }{ x }\]

OpenStudy (jiteshmeghwal9):

\[\int\limits \frac{2}{x(x^2+1)}dx\]\[ \int\limits \frac{2dx}{x^3(1+\frac{1}{x^2})}\]now substitute \(\Large \frac{1}{x^2}+1=u\) then \(\Large \frac{-2}{x^3}dx=du\)\[\int\limits \frac{-du}{u}\]replacing u again\[-\ln|u|\]\[-\ln \left( \frac{1}{x^2}+1 \right)\]

OpenStudy (jiteshmeghwal9):

\[\Large \text{I.F.}=e^{\int \frac{2}{x(x^2+1)}dx}=e^{-\ln \left(1+\frac{1}{x^2}\right)}\]

OpenStudy (caerus):

(1/x^2 +1)^-1

OpenStudy (jiteshmeghwal9):

yeah it is the integrating factor

OpenStudy (caerus):

can u help me with the right side again ^.^

OpenStudy (caerus):

y(1/x^2 +1)^-1=x + 1/x +c

OpenStudy (caerus):

\[\int\limits_{}^{}\frac{ (x^2+1)^2 }{ x }(1/x^2 +1)^-1\]

OpenStudy (caerus):

\[\int\limits_{}^{}\frac{ (x^2+1) }{ x }\]

OpenStudy (caerus):

@abb0t

OpenStudy (caerus):

\[ \frac{ y }{ (x^2+1)}=\frac{ (x+1)^2 }{ 2} +C\]

OpenStudy (caerus):

still end up the wrong answer

OpenStudy (caerus):

answer must be \[x^2y=\frac{ 1 }{ 4 }(x^2+1)^3+c(x^2+1)\]

OpenStudy (kainui):

Not really sure what's going on, but most people forget to slide their constant in the right place in the end, what I mean is you have something like this: \[(\mu y)' = \mu r(x)\] then \[\mu y = \int \mu r(x) dx\] then they'll usually write \[y = \frac{1}{\mu} \int \mu r(x) dx\] but they'll forget that when they integrate, their constant comes up here like this: \[y = \frac{1}{\mu} \int \mu r(x) dx + C\] when really it should be here like this since it distributes to both terms. \[y = \frac{1}{\mu} \int \mu r(x) dx + \frac{C}{\mu}\] Dunno if that helps or not, just a common mistake

OpenStudy (caerus):

anyone can point my mistake here ^.^ ... yeah @Kainui but i can see that error here..

OpenStudy (caerus):

cant*

OpenStudy (kainui):

Dunno, just start over from scratch maybe, it's good for ya. I'm working it out from start to finish with an integrating factor right now and I'll post the results in a sec.

OpenStudy (caerus):

yup @jiteshmeghwal9 , i continued it and i got an error, i dont know hwere it is. lol

OpenStudy (jiteshmeghwal9):

\[\int \frac{(x^2+1)^3}{x(x^2+1)}*(1+\frac{1}{x^2})^{-1}+k\]

OpenStudy (jiteshmeghwal9):

\[\int \frac{(x^2+1)^3}{x(x^2+1)}*\frac{x^2}{x^2+1} +k\]

OpenStudy (caerus):

\[\int\limits_{}^{}\frac{ (x^2+1) }{ x }\]

OpenStudy (caerus):

+k

OpenStudy (jiteshmeghwal9):

\[\int x(x^2+1)dx +k\]

OpenStudy (jiteshmeghwal9):

Solve this and u will get RHS

OpenStudy (kainui):

\[x(x^2+1)y'+2y=(x^2+1)^3\] Multiply by an integrating factor, it's completely arbitrary, whatever helps to solve your problem. \[\mu x(x^2+1)y'+2\mu y=\mu (x^2+1)^3\] Now let's go ahead and make up a differential equation to define our integrating function, \[[\mu x(x^2+1)]' = 2\mu\] The point being just to undo a power rule, go ahead and work out the power rule from my backwards step here, this is the whole point. \[(\mu x(x^2+1)y)' = \mu (x^2+1)^3\] Now we can integrate, although we still have to solve for \(\mu\) to finish. \[\mu x(x^2+1)y = \int \mu (x^2+1)^3dx\] \[y = \frac{1}{\mu x(x^2+1)} \int \mu (x^2+1)^3dx\] Alright time to solve our made up equation now. \[[\mu x(x^2+1)]' = 2\mu\] \[\mu' x(x^2+1) + \mu [x(x^2+1)]' = 2\mu\] \[\frac{\mu'}{\mu} =\frac{2 - [x(x^2+1)]'}{x(x^2+1)} \] \[\int \frac{d\mu}{\mu} = \int \frac{2 dx}{x(x^2+1)} -\int \frac{[x(x^2+1)]'}{x(x^2+1)}dx \] \[\ln \mu = \int \frac{2 dx}{x(x^2+1)} - \ln[x(x^2+1)]\] I don't know what that middle integral will be off the top of my head so I'll just stop here, the rest should follow though pretty easily since it's already all laid out. GL

OpenStudy (kainui):

haha I don't know if that helps or not to be honest, just did as much as was fun to me. :P

OpenStudy (jiteshmeghwal9):

That is a massive work lol

OpenStudy (caerus):

@jiteshmeghwal9 do you have time to help me again? lol

OpenStudy (jiteshmeghwal9):

\[(1+\frac{1}{x^2})^{-1}\]\[(\frac{x^2+1}{x^2})^{-1}\]\[(\frac{x^2}{x^2+1})\]

OpenStudy (jiteshmeghwal9):

Got it ?

OpenStudy (caerus):

yip, thanks, moving on to next problem ^.^